OpenStudy (anonymous):

Give the radius and the center for the following equation. x^2-10x+y^2-10y=-1 center:(___,___) radius=____

OpenStudy (anonymous):

the equation has tobe written in standard form first by completing the square \[(x-5)^{2}+(y-5)^{2}=-1+5^{2}\] \[(x-5)^{2}+(y-5)^{2}=24\] so \[Centre(5,5)\] \[r ^{2}=25\] \[r=\sqrt{24}=2\sqrt{6}\]

OpenStudy (anonymous):

But then if it`s 2 positive 5`s then it would be positive 10, not negative 10...

OpenStudy (anonymous):

\[x ^{2}-10x+(-5)^{2}-(-5)^{2}+y ^{2}-10y+(-y)^{2}-(-y)^{2}=-1\] \[x ^{2}-10x+(-5)^{2}+y ^{2}-10y+(-5)^{2}=-1+(-5)^{2}+(-5)^{2}\] \[(x-5)^{2}+(y-5)^{2}=-1+25+25=49\] centre always changes sign so C(5,5) sorry about th radius \[r=\sqrt{49}=7\]

OpenStudy (anonymous):

there we go, thanks :)